OpenStudy (anonymous):

would 1 - 2cos^2(y) = 2sin^2(y) ?????

OpenStudy (mathmate):

2 - 2cos^2(y) = 2sin^2(y)

OpenStudy (anonymous):

learning trig for the first time?

OpenStudy (anonymous):

^^

OpenStudy (anonymous):

ITS NOT

OpenStudy (anonymous):

Sorry. sin^2x + cos^2x = 1. So, 1-2cos^2x can't equal 2sin^2x.

OpenStudy (anonymous):

nope, just got confused. Butt: wouldn't the example which I gave form the equation that: 2sin^2x + 2cos^2x = 1, therefore sin^2x + cos^2x = 1/2 ????

OpenStudy (anonymous):

\( 1 - 2 \cos^2(y) = - \cos 2y \)

OpenStudy (anonymous):

thats equal to 2

OpenStudy (anonymous):

shinigami1m - sorry. I mis-read it first as 2 - 2cos^2x = 2sin^2x.

OpenStudy (anonymous):

yeahhh, that's what I was confused about. So help me solve this: 2cs^2y - siny - 1 = 0.

OpenStudy (anonymous):

cos or csc?

OpenStudy (anonymous):

cos^2y = 1-sin^2y. Substitute. You get a quadratic equation in siny. Then solve.

OpenStudy (anonymous):

sorry, cos, not cosec. ^Okayy, so lemme try that and see.

OpenStudy (amistre64):

s^2 + c^2 = 1 2s^2 + 2c^2 = 2 is what im thinking

OpenStudy (amistre64):

gotta use rads ;)

OpenStudy (anonymous):

WHAT IS THIS :O

OpenStudy (anonymous):

YOU FACE, SHAL'S. Finally rediscovered, after what: a million years??? :P

OpenStudy (anonymous):

I think this method might work: 1 - 2 cos square y = 2 sin square y(Given) Or, 1 = 2 sin square y + 2 cos square y Or, 1 = 2(sin square y + cos square y) We get, 1 = 2 (Impossible!!!!)

OpenStudy (anonymous):

Sometimes the question may be wrong.